# LIGHTS. A robust technique to identify galaxy edges

### v1

@@ 4.4 One dimensional edge detection:
In the top right panel of Fig. 5, we show the 2D map of the second derivative, representing changes along the angular direction and highlighting the natural rise of the edge of the galaxy. The location of the edge is also shown as a black dashed vertical line Fig. 4. Visual inspection tends to place the edge at a smaller radius, especially in profiles with smoother transitions between the inner and outer exponential regions (i.e., a lower parameter in Eq. 1). This is a perceptual bias: smoother transitions can mislead the eye into identifying a drop earlier than it actually occurs. In contrast, our edge detection method uses the second derivative to pinpoint the inflection point (see Appendix G for a visual example of this effect).
4.5 Bi-dimensional edge detection:
In order to characterize the edge of the galaxy in two dimensions we consider the following. To estimate the uncertainty in the intensity profile at each radial location (), we compute the standard deviation of the counts within an elliptical aperture . Under Gaussian approximation, the uncertainty in the total intensity measurement is given by . Therefore, dividing the elliptical aperture into 60-degree wedges reduces the number of pixels per wedge by a factor of 6 and increases the uncertainty by a factor of 2.5.
This division into 60 degree wedges still provides sufficient S/N for reliable edge detection. Therefore, to map the edge at different angular positions (), we divide the galaxy into 60-degree wedges. This means that at the semi-major axis, we define a wedge spanning PA = -30 deg to 30 deg, then shift by 30 degrees to define another wedge from PA = 0 deg to 60 deg, and so on. This means 12 wedges in total.
For each wedge, we extract the radial profile using masked data and sample it using the task astscript-radial-profile. As in the case of the 1D approach, the final value of is the mean between the minimum obtained with the numerical derivation and the analytical derivation of the polynomial fit. are represented as yellow dots in the upper panels of Fig. 5. The errors associated with these estimates corresponds to 5 (half width of the step). Fig. 6 shows the distribution of the and corresponding () compared to the value obtained with the one dimensional analysis. 1D and 2D methods yield consistent edge values, with angular positions of clustering around (few kpc), capturing small asymmetries in the disc’s structure. We calculate the average absolute deviation of all from , and express this value as a percentage of . This gives a mean asymmetry of for NGC 3486. The stellar mass densities at the edge cluster around the mean edge density , ranging between 0.75 and 4.5 . In Fig. 6, values are color-coded by their corresponding angular coordinate , revealing that neighboring angular sectors exhibit similar , suggesting local coherence in edge structure.
@@ 5 Discussion
Our results indicate that it is possible to accurately detect the edge of galaxies using an unbiased mathematical definition, using the stellar surface mass density distribution. NGC 3486 shows an almost symmetric disc (asymmetry ¡ 5) with a sharp decline beyond = 205 5 equal to 13.8 0.6 kpc at a distance of 13.6 Mpc. This corresponds to = 25.60 0.01 mag/arcsec2, = 25.95 0.01 mag/arcsec2, a mean surface stellar mass density of 1.3 0.1 and values between 0.75 to 4.5 .
@@
The results of this work therefore align with those of Chamba et al. (2022) for galaxies of similar mass and redshift ( ). They report values ranging from 6 to 20 kpc ( 16 kpc on average) and a stellar mass density at the edge for S0/a-Irr galaxies between 0.2 and 4 , with an average of 1.15 (see their Fig. 5). The galaxies in Chamba et al. (2022) that have both a similar stellar mass and an average color to that of NGC 3486 ( 0.3), have average edge radii greater than 10 kpc and tend to be larger than their redder counterparts.
5.1 What is driving the formation of a sharp truncation in NGC 3486?
@@
NGC 3486 shows a broken exponential stellar surface mass density profile, characterised by a change in slope that separates the decline of the inner region from the steeper outer regions. This transition appears sharp, with a ratio of 5 after removal of the effect of the PSF. This sharp transition is similar to that observed in the edge-on galaxies NGC 5907 and NGC 4565 (Martínez-Lombilla et al. 2018). The Near-Ultraviolet (NUV) and Far-Ultraviolet (FUV) profiles (see Fig. 8) also show a sharp drop, probably indicating negligible star formation beyond the edge. A similar break is also suggested in the infrared (IR) profile from Richards et al. (2016), where a change in slope is observed at 200 (see their Fig. A.1).
@@
The presence of this clear edge suggests an ongoing star formation process. Over time, this sharp boundary is expected to fade as stellar migration redistributes stars from the inner regions inside into the outer disc (Sellwood 2014; Debattista et al. 2017), gradually erasing signatures of past star formation. Further evidence for this star formation link can be found in the color profile. Within the galaxy shows active star formation, while beyond this radius the color becomes redder, possibly due to stellar migration. This is consistent with the results of Martínez-Lombilla et al. (2018), who interpret the gradual reddening beyond the edge as passive aging of the stellar population.
@@
Under the hypothesis that marks the limit of in-situ star formation, the formation and evolution of the edge is intimately linked to the abundance of HI gas. The absence of star formation in the outer disc is likely tied to the availability of cold gas and feedback effects. Observations of HI and molecular gas distributions in galaxies with similar profiles show that gas densities drop below the critical threshold for star formation beyond the edge. Using Very Large Array (VLA) data, Richards et al. (2016) found that NGC 3486 has an asymmetric HI gas distribution that extends from the northwest and southeast sides of the HI disc. They also infer a mean inclination from the HI data at 200 of deg that is consistent with the axis ratio of 0.765 found in this work. In Fig. 14 (see Appendix H) we have tentatively superimposed the HI contours from Richards et al. (2016) on the LIGHTS data. The asymmetric HI gas distribution is further analyzed by Smith et al. (2021), who reports two HI arcs extending from the southeastern and northwestern parts of NGC 3486 (see their Fig. A.9) and rotating with the rest of the galaxy. We hypothesize that the observed HI asymmetry could be due to an interaction with a satellite galaxy (PGC 033184; see Fig. 9) located to the north of NGC 3486 at z = 0.008. The gas content of NGC 3486 was also studied by Broeils & Van Woerden (1994) using short 21-cm line observations from the Westerbork Synthesis Radio Telescope. Fig. 7 shows a comparison between their derived surface gas mass density profile (blue) and the surface stellar mass density of this work (orange). We have adapted their profile to the distance assumed in this paper. At , the surface HI gas mass density reaches , consistent with observations of nearby spiral galaxies (Martin & Kennicutt Jr 2001; Leroy et al. 2008). In NGC 3486, the star formation efficiency at is around 25, as the surface gas mass density at this radius is 3.8 times higher than the stellar mass density. This result is therefore consistent with theoretical models suggesting that the star formation threshold in galaxies is regulated by thermal instabilities arising from the transition between warm and cold gas phases (Schaye 2004).
@@ 5.2 Method Limitations and future directions
We evaluate the performance and implications of our new edge detection method, which mitigates the effects of scattered light and help in identifying the edge location (and therefore the characterization of the size using this feature as a proxy) in galaxies. This method is crucial in deep imaging, providing a robust approach to identifying galaxy edges based on second derivatives. This technique also provides the first wedge-based characterization of the border of galaxies. The code is modular and automated, including profile extraction, derivative calculation, and edge detection in both 1D and 2D. We acknowledge that masking low surface brightness features, deciding on the sampling of the surface brightness and mass profiles, or adapting wedge widths still involve manual decisions and may carry subjective elements. Computing the second derivative can introduce numerical errors, which may stem from point spacing or limited floating-point precision. Our approach minimizes smoothing to preserve data authenticity, but this remains a balancing challenge. While computationally efficient overall, the most resource-intensive step remains the PSF correction of the data. Future work (Golini et al., in prep.) will explore the method’s reproducibility across datasets with varying depths, noise levels, resolution (i.e., different sampling of the mass profile) and PSF models, and expand its application to a broader range of galaxy properties (e.g., inclination, size, mass).
@@ 6 Conclusions
We have introduced a robust method for measuring the edges of galaxies by leveraging the second derivative of the stellar surface mass density profile. This approach serves as a complementary quantitative alternative to traditional visual-based methods (Martínez-Lombilla et al. 2018; Trujillo et al. 2020; Díaz-García et al. 2022; Chamba et al. 2022; Buitrago & Trujillo 2024) for identifying galaxy boundaries. Our technique enables both one- and two-dimensional analysis, allowing for a comprehensive characterization of galaxy outskirts, asymmetries, and structural evolution. By applying our method to deep imaging observations of NGC 3486 from the LIGHTS (Trujillo et al. 2021; Zaritsky et al. 2024) survey ( = mag/arcsec2 and = mag/arcsec2; 3 in 10 10 boxes), we identified an outer edge at 205 5, in strong agreement with the visual boundary. Our wedge-based analysis quantified the galaxy’s asymmetry at 5, confirming the consistency between the isophotal ellipse parameters and the edge contours at different angular directions. The sharp transition at the edge occurs at a stellar surface mass density of /pc2, reinforcing the possible link between galaxy edges and the threshold of in-situ star formation (Quirk 1972; Kennicutt 1989; Schaye 2004; Roškar et al. 2008).
The computational efficiency and adaptability of this method make it well-suited for integration into survey pipelines, ensuring consistent edge detection across different datasets, such as LSST (Ivezić et al. 2019), Euclid (Scaramella et al. 2022), Roman (Bailey et al. 2023) and ARRAKIHS (Guzmán 2024). In summary, this method offers a promising advance in galaxy edge detection, while aligning well with the needs of deep imaging surveys (such as background and scattered light removal). Continued development will enable its application across varied datasets, paving the way for improved understanding of galaxy structures and the processes that shape them.
@@ Acknowledgements.
- Díaz-García et al. (2022) Díaz-García, S., Comerón, S., Courteau, S., et al. 2022, A&A, 667, A109
@@
- Martín-Navarro et al. (2012) Martín-Navarro, I., Bakos, J., Trujillo, I., et al. 2012, MNRAS, 427, 1102
- Martínez-Lombilla et al. (2023) Martínez-Lombilla, C., Infante-Sainz, R., Jiménez-Ibarra, F., et al. 2023, A&A, 678, A62
- Martínez-Lombilla & Knapen (2019) Martínez-Lombilla, C. & Knapen, J. H. 2019, A&A, 629, A12
@@ Appendix B Color image of LIGHTS data
Fig. 9 shows the LIGHTS data of NGC 3486. The color image with gray-scaled background is constructed using astscript-color-faint-gray (Infante-Sainz & Akhlaghi 2024), by combining the and filters. The black and white background is from the band.
@@ Appendix D Structural parameters of Model1 and Model2
The IMFIT parameters of the simulated Model1 and Model2 share similarities but differ in key aspects. To create both models we use a BrokenExponential profile with a position angle of -20 degrees (clockwise from the X axis), an inclination of 39.6 degrees and an inner scale length of 75 pixels. The location of the transition between inner and outer exponential declines is set to 200 pixels, and is equal to 10 in both models. Differences arise in the outer scale length , where in Model1 it is set to 30 pixels, while in Model2 to 25 pixels. Additionally, in the halo Exponential component, Model2 has central intensity value of 0.0004 counts/pix and scale length of 270 pixels.
@@ Appendix E Characterising the Position Angle and axis ratio of NGC3486
To extract the surface brightness profile of NGC 3486, we first use the ellipse routine in photutils (Jedrzejewski 1987; Bradley et al. 2019) to fit elliptical isophotes to the combined and band data. This method provides a representative PA and ellipticity of the outer galaxy regions. In the initial run of ellipse, all parameters vary freely. In the subsequent run, we fix the center to the median coordinates derived in the first pass (R.A. (J2000) = 165.0994 deg, Dec. (J2000) = 28.9751 deg)999The uncertainty associated with the selection of the center of the galaxy is 0.3 in R.A. and Dec (Jedrzejewski 1987).. The fitting begins at 2 with linearly spaced elliptical annuli. We find an average AR of and PA of (clockwise from the X-axis) between radii of 170 and 220. These values align with prior measurements, such as AR = 0.792 and PA = (clockwise from the X-axis; Smith et al. 2021) and PA = (clockwise from the X-axis) with an inclination of 36.8 from gas kinematics (Richards et al. 2016).
@@ Appendix F Tests on Wiener deconvolution
PSF, saturation and noise. To evaluate the accuracy of the Wiener-Hunt deconvolution method, we applied it to the simulated disc Model1. By comparing the deconvolved images to their original (not convolved) counterparts, we assess how well the method retrieves the intrinsic averaged surface brightness profiles and we visually compare the background variations. The profiles are extracted using fixed ellipses for each model, and our findings are discussed below.
Fig. 11 presents a comparative analysis of disc models, illustrating the effects of the PSF and Wiener-Hunt deconvolution under different conditions. The first column represents the base model, while subsequent columns introduce noise and saturation effects. Each row shows a different processing stage. The first row displays the original model (assuming infinite S/N), the same model with added Gaussian ( = 0.003) noise using the task mknoise-sigma-from-mean function in Gnuastro (Noisy model), the Noisy model with inner 20 replaced with Not a Number (NaNs) simulating saturation (Sat. Noisy model), and the original model adding higher Gaussian ( = 0.03) noise (Noisy2 model). The Noisy model simulates an image with a limiting surface brightness of 29.1 mag/arcsec2 (3; 10 10), while the Noisy2 model corresponds to 26.6 mag/arcsec2 (3; 10 10). These values are representative of typical surveys such as SDSS and Stripe 82.
In the second row, the same models are convolved with the LIGHTS’ PSF (Sedighi et al. 2025) using the --psf option in IMFIT.
The PSF decline as power law () with an exponent = -2.1.
@@
Comparison with wavelet-regularized deconvolution. Wiener-Hunt deconvolution is not the only method we can use to address the PSF-Stellar halo degeneracy. We compare Wiener-Hunt deconvolution with wavelet-regularized deconvolution techniques. Before delving into this comparison, we introduce the concept of wavelets as follows. The quality of the deconvolved images can be improved by performing a regularized deconvolution. If the noise term in the image formation assumption of Eq. 2 is Gaussian with zero mean and diagonal covariance matrix, the deconvolution can be performed in a Bayesian framework by optimizing the following metric:
@@
where represents the isotropic undecimated wavelet transform (IUWT), also known as the starlet transform (Starck & Murtagh 1994; Starck et al. 2006). The IUWT is a redundant wavelet transform that decomposes an image into multiple frequency components, providing a sparse representation well-suited for astronomical images. For a more detailed explanation of the mathematical foundations behind wavelet-regularized deconvolution methods, we refer the reader to the following works: Starck & Murtagh (1994, 2007); Starck et al. (2010); Carrillo et al. (2012). Fig. 12 shows a cropped region of the outskirts of NGC 3486 (3′ 3′) before and after deconvolution and compares the averaged surface brightness profiles obtained with Wiener-Hunt and Wavelet deconvolution methods using fixed elliptical apertures. We find close agreement up to 280 corresponding to = 29 mag/arcsec2.

### v2

@@ 4.4 One-dimensional edge detection:
In the top right panel of Fig. 5, we show the 2D map of the second derivative, representing changes in the angular direction and highlighting the natural rise of the edge of the galaxy. The location of the edge is also shown as a vertical dashed black line in Fig. 4. Visual inspection tends to place the edge at a smaller radius, especially in profiles with smoother transitions between the inner and outer exponential regions (i.e., a lower parameter in Eq. 1). This is a perceptual bias: smoother transitions can mislead the eye into identifying a drop earlier than it actually occurs. In contrast, our edge detection method uses the second derivative to pinpoint the inflection point (see Appendix G for a visual example of this effect).
4.5 Bidimensional edge detection:
In order to characterize the edge of the galaxy in two dimensions, we considered the following. To estimate the uncertainty in the intensity profile at each radial location (), we computed the standard deviation of the counts within an elliptical aperture . Under Gaussian approximation, the uncertainty in the total intensity measurement is given by . Therefore, dividing the elliptical aperture into 60-degree wedges reduces the number of pixels per wedge by a factor of 6 and increases the uncertainty by a factor of 2.5.
This division into 60 degree wedges still provides a sufficient S/N for reliable edge detection. Therefore, to map the edge at different angular positions (), we divided the galaxy into 60-degree wedges. This means that at the semimajor axis, we defined a wedge spanning PA = -30 deg to 30 deg, then shifted by 30 degrees to define another wedge from PA = 0 deg to 60 deg, and so on. This means 12 wedges in total.
For each wedge, we extracted the radial profile using masked data and sampled it using the task astscript-radial-profile. As in the case of the 1D approach, the final value of is the mean between the minimum obtained with the numerical derivation and the analytical derivation of the polynomial fit. are represented as yellow dots in the upper panels of Fig. 5. The errors associated with these estimates correspond to 5 (the half width of the step). Fig. 6 shows the distribution of the and corresponding () compared to the value obtained with the 1D analysis. The 1D and 2D methods yield consistent edge values, with angular positions of clustering around (few kiloparsec), capturing small asymmetries in the disk’s structure. We calculated the average absolute deviation of all from , and express this value as a percentage of . This gives a mean asymmetry of for NGC 3486. The stellar mass densities at the edge cluster around the mean edge density , ranging between 0.75 and 4.5 . In Fig. 6, values are color-coded by their corresponding angular coordinate, , revealing that neighboring angular sectors exhibit similar , suggesting local coherence in edge structure.
@@ 5 Discussion
Our results indicate that it is possible to accurately detect the edge of galaxies through an unbiased mathematical definition, using the stellar surface mass density distribution. NGC 3486 shows an almost symmetric disk (asymmetry ¡ 5) with a sharp decline beyond = 205 5 equal to 13.8 0.6 kpc at a distance of 13.6 Mpc. This corresponds to = 25.60 0.01 mag/arcsec2, = 25.95 0.01 mag/arcsec2, a mean surface stellar mass density of 1.3 0.1 , and values between 0.75 to 4.5 .
@@
The results of this work therefore align with the ones of Chamba et al. (2022) for galaxies of a similar mass and redshift ( ). They report values ranging from 6 to 20 kpc ( 16 kpc on average) and a stellar mass density at the edge for S0/a-Irr galaxies between 0.2 and 4 , with an average of 1.15 (see their Fig. 5). The galaxies in Chamba et al. (2022) that have both a similar stellar mass and an average color to that of NGC 3486 ( 0.3) have average edge radii greater than 10 kpc and tend to be larger than their redder counterparts.
5.1 The formation of a sharp truncation in NGC 3486
@@
NGC 3486 shows a broken exponential stellar surface mass density profile, characterized by a change in slope that separates the decline of the inner region from the steeper outer regions. This transition appears sharp, with a ratio of 5 after removal of the effect of the PSF. This sharp transition is similar to that observed in the edge-on galaxies NGC 5907 and NGC 4565 (Martínez-Lombilla et al. 2018). The near-ultraviolet (NUV) and far-ultraviolet (FUV) profiles (see Fig. 8) also show a sharp drop, probably indicating negligible star formation beyond the edge. A similar break is also suggested in the infrared profile from Richards et al. (2016), where a change in slope is observed at 200 (see their Fig. A.1).
@@
The presence of this clear edge suggests an ongoing star formation process. Over time, this sharp boundary is expected to fade as stellar migration redistributes stars from the inner regions inside into the outer disk (Sellwood 2014; Debattista et al. 2017), gradually erasing signatures of past star formation. Further evidence for this star formation link can be found in the color profile. Within the galaxy shows active star formation, while beyond this radius the color becomes redder, possibly due to stellar migration. This is consistent with the results of Martínez-Lombilla et al. (2018), who interpret the gradual reddening beyond the edge as passive aging of the stellar population.
@@
Under the hypothesis that marks the limit of in situ star formation, the formation and evolution of the edge is intimately linked to the abundance of HI gas. The absence of star formation in the outer disk is likely tied to the availability of cold gas and feedback effects. Observations of HI and molecular gas distributions in galaxies with similar profiles show that gas densities drop below the critical threshold for star formation beyond the edge. Using Very Large Array (VLA) data, Richards et al. (2016) found that NGC 3486 has an asymmetric HI gas distribution that extends from the northwest and southeast sides of the HI disk. They also infer a mean inclination from the HI data at 200 of deg that is consistent with the AR of 0.765 found in this work. In Fig. 14 (see Appendix H) we have tentatively superimposed the HI contours from Richards et al. (2016) on the LIGHTS data. The asymmetric HI gas distribution is further analyzed by Smith et al. (2021), who reports two HI arcs extending from the southeastern and northwestern parts of NGC 3486 (see their Fig. A.9) and rotating with the rest of the galaxy. We hypothesize that the observed HI asymmetry could be due to an interaction with a satellite galaxy (PGC 033184; see Fig. 9) located to the north of NGC 3486 at z = 0.008. The gas content of NGC 3486 was also studied by Broeils & Van Woerden (1994) using short 21-cm line observations from the Westerbork Synthesis Radio Telescope. Fig. 7 shows a comparison between their derived surface gas mass density profile (blue) and the surface stellar mass density of this work (orange). We have adapted their profile to the distance assumed in this paper. At , the surface HI gas mass density reaches , consistent with observations of nearby spiral galaxies (Martin & Kennicutt Jr 2001; Leroy et al. 2008). In NGC 3486, the star formation efficiency at is around 25, as the surface gas mass density at this radius is 3.8 times higher than the stellar mass density. This result is therefore consistent with theoretical models suggesting that the star formation threshold in galaxies is regulated by thermal instabilities arising from the transition between warm and cold gas phases (Schaye 2004).
@@ 5.2 Method limitations and future directions
We evaluated the performance and implications of our new edge detection method, which mitigates the effects of scattered light and helps to identify the edge location (and therefore the characterization of the size using this feature as a proxy) in galaxies. This method is crucial in deep imaging, providing a robust approach to identifying galaxy edges based on second derivatives. This technique also provides the first wedge-based characterization of the border of galaxies. The code is modular and automated, including profile extraction, derivative calculation, and edge detection in both 1D and 2D. We acknowledge that masking LSB features, deciding on the sampling of the surface brightness and mass profiles, or adapting wedge widths still involves manual decisions and may carry subjective elements. Computing the second derivative can introduce numerical errors, which may stem from point spacing or limited floating-point precision. Our approach minimizes smoothing to preserve data authenticity, but this remains a balancing challenge. While computationally efficient overall, the most resource-intensive step remains the PSF correction of the data. Future work (Golini et al., in prep.) will explore the method’s reproducibility across datasets with varying depths, noise levels, resolution (i.e., different sampling of the mass profile), and PSF models, and expand its application to a broader range of galaxy properties (e.g., inclination, size, mass).
@@ 6 Conclusions
We have introduced a robust method of measuring the edges of galaxies by leveraging the second derivative of the stellar surface mass density profile. This approach serves as a complementary quantitative alternative to traditional visual-based methods (Martínez-Lombilla et al. 2018; Trujillo et al. 2020; D´ıaz-Garc´ıa et al. 2022; Chamba et al. 2022; Buitrago & Trujillo 2024) of identifying galaxy boundaries. Our technique enables both 1D and 2D analysis, allowing for a comprehensive characterization of galaxy outskirts, asymmetries, and structural evolution. By applying our method to deep imaging observations of NGC 3486 from the LIGHTS (Trujillo et al. 2021; Zaritsky et al. 2024) survey ( = mag/arcsec2 and = mag/arcsec2; 3 in 10 10 boxes), we identified an outer edge at 205 5, in strong agreement with the visual boundary. Our wedge-based analysis quantified the galaxy’s asymmetry at 5, confirming the consistency between the isophotal ellipse parameters and the edge contours at different angular directions. The sharp transition at the edge occurs at a stellar surface mass density of /pc2, reinforcing the possible link between galaxy edges and the threshold of in situ star formation (Quirk 1972; Kennicutt 1989; Schaye 2004; Roškar et al. 2008).
The computational efficiency and adaptability of this method make it well suited for integration into survey pipelines, ensuring consistent edge detection across different datasets, such as LSST (Ivezić et al. 2019), Euclid (Scaramella et al. 2022), Roman (Bailey et al. 2023), and ARRAKIHS (Guzmán 2024). In summary, this method offers a promising advance in galaxy edge detection, while aligning well with the needs of deep imaging surveys (such as background and scattered light removal). Continued development will enable its application across varied datasets, paving the way for an improved understanding of galaxy structures and the processes that shape them.
@@ Acknowledgements.
- D´ıaz-Garc´ıa et al. (2022) Díaz-García, S., Comerón, S., Courteau, S., et al. 2022, A&A, 667, A109
@@
- Mart´ın-Navarro et al. (2012) Martín-Navarro, I., Bakos, J., Trujillo, I., et al. 2012, MNRAS, 427, 1102
- Mart´ınez-Lombilla et al. (2023) Martínez-Lombilla, C., Infante-Sainz, R., Jiménez-Ibarra, F., et al. 2023, A&A, 678, A62
- Mart´ınez-Lombilla & Knapen (2019) Martínez-Lombilla, C. & Knapen, J. H. 2019, A&A, 629, A12
@@ Appendix B Color image of LIGHTS data
Fig. 9 shows the LIGHTS data of NGC 3486. The color image with gray-scaled background was constructed using astscript-color-faint-gray (Infante-Sainz & Akhlaghi 2024), by combining the and filters. The black and white background is from the band.
@@ Appendix D Structural parameters of Model1 and Model2
The IMFIT parameters of the simulated Model1 and Model2 share similarities but differ in key aspects. To create both models we used a BrokenExponential profile with a PA of -20 degrees (clockwise from the X axis), an inclination of 39.6 degrees and an inner scale length of 75 pixels. The location of the transition between inner and outer exponential declines was set to 200 pixels, and is equal to 10 in both models. Differences arise in the outer scale length , where in Model1 it is set to 30 pixels, while in Model2 to 25 pixels. Additionally, in the halo Exponential component, Model2 has central intensity value of 0.0004 counts/pix and scale length of 270 pixels.
@@ Appendix E Characterising the position angle and axis ratio of NGC3486
To extract the surface brightness profile of NGC 3486, we first used the ellipse routine in photutils (Jedrzejewski 1987; Bradley et al. 2019) to fit elliptical isophotes to the combined and band data. This method provides a representative PA and ellipticity of the outer galaxy regions. In the initial run of ellipse, all parameters varied freely. In the subsequent run, we fixed the center to the median coordinates derived in the first pass (R.A. (J2000) = 165.0994 deg, Dec. (J2000) = 28.9751 deg)999The uncertainty associated with the selection of the center of the galaxy is 0.3 in R.A. and Dec (Jedrzejewski 1987).. The fitting began at 2 with linearly spaced elliptical annuli. We found an average AR of and PA of (clockwise from the X axis) between radii of 170 and 220. These values align with prior measurements, such as AR = 0.792 and PA = (clockwise from the X axis; Smith et al. 2021) and PA = (clockwise from the X axis) with an inclination of 36.8 from gas kinematics (Richards et al. 2016).
@@ Appendix F Tests on Wiener deconvolution
PSF, saturation and noise. To evaluate the accuracy of the Wiener-Hunt deconvolution method, we applied it to the simulated disk Model1. By comparing the deconvolved images to their original (not convolved) counterparts, we assessed how well the method retrieves the intrinsic averaged surface brightness profiles and we visually compared the background variations. The profiles were extracted using fixed ellipses for each model. Our findings are discussed below.
Fig. 11 presents a comparative analysis of disk models, illustrating the effects of the PSF and Wiener-Hunt deconvolution under different conditions. The first column represents the base model, while subsequent columns introduce noise and saturation effects. Each row shows a different processing stage. The first row displays the original model (assuming infinite S/N), the same model with added Gaussian ( = 0.003) noise using the task mknoise-sigma-from-mean function in Gnuastro (Noisy model), the Noisy model with inner 20 replaced with Not a Number (NaNs) simulating saturation (Sat. Noisy model), and the original model adding higher Gaussian ( = 0.03) noise (Noisy2 model). The Noisy model simulates an image with a limiting surface brightness of 29.1 mag/arcsec2 (3; 10 10), while the Noisy2 model corresponds to 26.6 mag/arcsec2 (3; 10 10). These values are representative of typical surveys such as Stripe 82 and SDSS.
In the second row, the same models were convolved with the LIGHTS’ PSF (Sedighi et al. 2025) using the --psf option in IMFIT.
The PSF declines as a power law () with an exponent = -2.1.
@@
Comparison with wavelet-regularized deconvolution. Wiener-Hunt deconvolution is not the only method we can use to address the PSF-Stellar halo degeneracy. We compared Wiener-Hunt deconvolution with wavelet-regularized deconvolution techniques. Before delving into this comparison, we introduce the concept of wavelets as follows. The quality of the deconvolved images can be improved by performing a regularized deconvolution. If the noise term in the image formation assumption of Eq. 2 is Gaussian with zero mean and diagonal covariance matrix, the deconvolution can be performed in a Bayesian framework by optimizing the following metric:
@@
where represents the isotropic undecimated wavelet transform (IUWT), also known as the starlet transform (Starck & Murtagh 1994; Starck et al. 2006). The IUWT is a redundant wavelet transform that decomposes an image into multiple frequency components, providing a sparse representation well-suited for astronomical images. For a more detailed explanation of the mathematical foundations behind wavelet-regularized deconvolution methods, we refer the reader to the following works: Starck & Murtagh (1994, 2007); Starck et al. (2010); Carrillo et al. (2012). Fig. 12 shows a cropped region of the outskirts of NGC 3486 (3′ 3′) before and after deconvolution and compares the averaged surface brightness profiles obtained with Wiener-Hunt and wavelet-regularized deconvolution methods using fixed elliptical apertures. We find close agreement up to 280 corresponding to = 29 mag/arcsec2.